A Bibliography of Publications of Juliane Dehnert

Juliane Dehnert Technical University Berlin Fachbereich 13, Informatik

Institute for Computation and Information Structures Sekr. E-N 7 Einsteinufer 17 D-10587 Berlin

Germany

Tel: ?n/a? FAX: ?n/a?

E-mail: dehnert@cs.tu-berlin.de (Internet)

13 October 2017 Version 1.06

Abstract

This bibliography records publications of Juliane Dehnert.

Title word cross-reference

13th [DGN01].

Advanced [DGN01, L^{+}01]. Applications [L^{+}01]. April [L^{+}01]. Atomicity [DDGJ01, DDGJ00].

Beijing [L^{+}01]. Business [Deh01b, Deh01a, DR01].

CAiSE [DGN01]. China [L^{+}01]. choice [Deh02]. CODAS [L^{+}01]. Colored [DFZ00b]. conference [DGN01].

Controllability [Deh01b, Deh02]. controllable [Deh02]. Cooperative [L^{+}01]. Customized [DDGJ00, DDGJ01].

Database [L^{+}01].

engineering [DGN01]. Evaluation [DFZ00a, DFZ00b]. Expressing [Deh01b, Deh02].

Four [Deh01a].

information [DGN01]. Interlaken [DGN01]. International [L^{+}01, DGN01].

June [DGN01].

Modeling [DFZ00a, DFZ00b, ED03]. Models [Deh01a].

Nets [DFZ00b, ED03]. Non [Deh02]. Non-controllable [Deh02].

Performance [DFZ00a, DFZ00b]. Petri [DFZ00b, ED03]. proceedings [DGN01, L^{+}01]. Process [Deh01a]. Processes [Deh01b, DR01, Deh02].

Reactive [ED03]. Relaxed [DR01]. robustness [Deh02].

Sound [Deh01a]. Soundness [DR01]. Specification [DDGJ01, DDGJ00]. Steps [Deh01a]. Stochastic [DFZ00b]. Switzerland [DGN01]. Symposium [L^{+}01]. Systematic [Deh01a]. Systems [DFZ00a, L^{+}01, DGN01].

Third [L^{+}01]. Transactional [DDGJ01, DDGJ00].

Workflow [DFZ00a, DFZ00b, DDGJ01, Deh02, DDGJ00, ED03].

References

Derks:2000:CAS

[DDGJ00] Wijnand Derks, Juliane Dehnert, Paul Grefen, and Willem Jonker. Customized atomicity specification for transactional workflow. CTIT Technical Report 00-24, University of Twente, Enschede, The Netherlands, 2000.

Derks:2001:CAS

[DDGJ01] Wijnand Derks, Juliane Dehnert, Paul Grefen, and Willem Jonker. Customized atomicity specification for transactional workflow. In Lu et al. [L^{+}01], pages 140– 147. ISBN 0-7695-1128-7, 0-7695-1129-5 (case), 0-7695-1130-9 (microfiche). LCCN ????

Dehnert:2001:FSS

[Deh01a] J. Dehnert. Four systematic steps towards sound business process models. In H. Weber, H. Ehrig, and W. Reisig, editors, 2nd Int. Colloquium on Petri Net Technologies for Modelling Communication Based Systems, pages 55– 64. Fraunhofer Gesellschaft ISST, ????, September 2001.

Dehnert:2001:ECB

[Deh01b] Juliane Dehnert. Expressing the controllability of business processes. Petri Net Newsletter, 61: 9–17, 2001. CODEN ???? ISSN ???? Column: Refereed Technical Contribution.

Dehnert:2002:NCC

[Deh02] Juliane Dehnert. Non-controllable choice robustness expressing the controllability of workflow processes. In Application and theory of Petri nets 2002, volume 2360 of Lecture Notes in Computer Science, pages 121– 141. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002.

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2360/23600121.htm;
 http://link.springer-ny.com/
link/service/series/0558/papers/
2360/23600121.pdf.

Dehnert:2000:MPE

[DFZ00a] J. Dehnert, A. Freiheit, and A. Zimmermann. Modeling and performance evaluation of workflow systems. In Proc. 4th. World Multiconference on Systemics, Cybernetics and Informatics (SCI’2000), Orlando, Florida, volume VIII, pages 632–637, 2000.

Dehnert:2000:WMP

[DFZ00b] J. Dehnert, A. Freiheit, and A. Zimmermann. Workflow modeling and performance evaluation with colored stochastic Petri nets. In Bringing Knowledge to Business Processes, Workshop in the American Association for Artificial Intelligence (AAAI) Spring Symposium Series 2000, pages 139–141, 2000.

Dittrich:2001:AIS

[DGN01] Klaus R. Dittrich, Andreas Geppert, and Moira C. Norrie, editors. Advanced information systems engineering: 13th international conference, CAiSE 2001, Interlaken, Switzerland, June 4– 8, 2001: proceedings, volume 2068 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42215-3 (paperback). LCCN QA76.758 .C33 2001.

URL http://link.springer-
ny.com/link/service/series/
0558/tocs/t2068.htm.

Dehnert:2001:RSB

[DR01] J. Dehnert and P. Rittgen. Relaxed soundness of business processes. In Dittrich et al. [DGN01], pages 157– 170. ISBN 3-540-42215-3 (paperback). LCCN QA76.758 .C33 2001.

URL http://link.springer-

                                        ny.com/link/service/series/
                                        0558/bibs/2068/20680157.htm;
                                         http://link.springer-ny.com/
                                        link/service/series/0558/papers/
                                        2068/20680157.pdf.

Eshuis:2003:RPN

[ED03] Rik Eshuis and Juliane Dehnert. Reactive Petri nets for workflow modeling. In Applications and theory of Petri nets 2003, volume 2679 of Lecture Notes in Computer Science, pages 296–315. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Lu:2001:PTI

[L^{+}01] Hongjun Lu et al., editors. The proceedings of the Third International Symposium on Cooperative Database Systems for Advanced Applications: CODAS 2001, Beijing, China, April 23–24, 2001. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1128-7, 0-7695-1129-5 (case), 0-7695-1130-9 (microfiche). LCCN ????